# Stochastic synaptic dynamics under learning

### v1

@@ IV.3 Forgetting on the macroscopic time scale
for which we do not know an explicit solution. Still, Eq. (46) can be evaluated numerically. $K^{(k)}$ and $G^{(k)}$ are shown and compared to simulations in Fig. 4(a,b); within a few hundert patterns, $K^{(k)}$ and $G^{(k)}$ decay to their equilibrium values $m_{0}$ and $V_{0}$, respectively.
Recall
@@
$\frac{f_{s}}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,M_{11}^{k}}{\sqrt{2f_{c}MV_{11}^{k}}}\right)+\frac{1-f_{s}}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,M_{01}^{k}}{\sqrt{2f_{c}MV_{01}^{k}}}\right)=f_{s},$ (52)
@@
$a_{k}^{\text{fp}}$ $=f_{s}^{-1}\int_{s_{\ast}}^{\infty}ds\,(1-f_{s})\mathcal{N}(s|f_{c}MM_{01}^{k},f_{c}MV_{01}^{k}).$
$=f_{s}^{-1}(1-f_{s})\frac{1}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,M_{01}^{k}}{\sqrt{2f_{c}MV_{01}^{k}}}\right).$ (53)
@@ Appendix A Numerical methods
Simulations of synapses and neurons are implemented in cython [63]. For the LIF neurons, Eq. (4), membrane voltages are integrated with the Euler-Maruyama scheme. In the network, Eq. (5) and Eq. (6), spike detection in one time step results in voltage reset and spike delivery in the same time step. The synaptic model is simulated by integrating Eq. (3) with the Euler scheme. Additionally, whenever the weight is set to a negative value, it is clipped to zero. For the experimentally inspired parameters used in the present paper, this almost never happens. For both Eq. (4) and Eq. (3), the Euler time step is $\Delta t=10^{-4}\tau_{m}$.
Eq. (24) and Eq. (32) are integrated numerically using a Runge-Kutta scheme of order 5(4).
@@ Appendix B Response functions of the LIF neuron
The susceptibilities required in Eq. (17) are the susceptibility of the LIF neuron Eq. (4) to mean- and noise-intensity-modulations, respectively. These have been derived using Fokker-Plank theory in [27],
@@
where $\mathcal{D}_{z}(x)$ is Whittaker’s parabolic cylinder function. For the drift coefficient we need to evaluate Eq. (54) and Eq. (55) at imaginary frequencies. This is convenient, as it only requires to evaluate $\mathcal{D}_{z}(x)$ at $z\in\mathbb{R}$. For the figures, we use the implementation $\mathtt{scipy.special.pbdv(z,x)}$.
@@ Appendix C Approximation of $D^{(2)}$
$\left\langle\dot{w}(t^{\prime})\dot{w}(t^{\prime\prime})\Big{|}w_{\text{traj}}(t)=w\right\rangle$ $=\nu r(w)\,\delta(t^{\prime}-t^{\prime\prime})\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)^{2}$ (61)
@@
$D^{(2)}(w)=\frac{1}{2}r(w)\nu\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)^{2}=\frac{1}{4}r(w)\nu(\Delta_{c}^{2}\tau_{c}+r_{ac}^{2}w^{2}\tau_{ac}).$ (62)
@@ Appendix D Mean-field theory of the recurrent network
Thus, the input to all neurons is determined by the firing rate $r_{1}$, $r_{0}$, and $r_{I}$, which in turn are determined by Eq. (13). This self-consistent set of equations is solved using a damped saddle-point iteration.

### v2

@@ IV.3.1 Trace degradation
for which we do not know an explicit solution. Still, Eq. (46) can be evaluated numerically. $K^{(k)}$ and $G^{(k)}$ are shown and compared to simulations in Fig. 4(a,b); within a few hundred patterns, $K^{(k)}$ and $G^{(k)}$ decay to their equilibrium values $m_{0}$ and $V_{0}$, respectively.
IV.3.2 Recall
@@
$\frac{f_{s}}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,K_{11}^{(k)}}{\sqrt{2f_{c}MG_{11}^{(k)}}}\right)+\frac{1-f_{s}}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,K_{01}^{(k)}}{\sqrt{2f_{c}MG_{01}^{(k)}}}\right)=f_{s},$ (52)
@@
$a_{k}^{\text{fp}}$ $=f_{s}^{-1}\int_{s_{\ast}}^{\infty}ds\,(1-f_{s})\mathcal{N}(s|f_{c}MK_{01}^{(k)},f_{c}MG_{01}^{(k)})$
$=f_{s}^{-1}(1-f_{s})\frac{1}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,K_{01}^{(k)}}{\sqrt{2f_{c}MG_{01}^{(k)}}}\right).$ (53)
@@ Appendix A Numerical methods
Simulations of synapses and neurons are implemented in cython [63]. For the LIF neurons, Eq. (4), membrane voltages are integrated with the Euler-Maruyama scheme. In the network, Eqs. (5) and (6), spike detection in one time step results in voltage reset and spike delivery in the same time step. The synaptic model is simulated by integrating Eq. (3) with the Euler scheme. Additionally, whenever the weight is set to a negative value, it is clipped to zero. For the experimentally inspired parameters used in the present paper, this almost never happens. For both Eq. (4) and Eq. (3), the Euler time step is $\Delta t=10^{-4}\tau_{m}$.
Equations (24) and (32) are integrated numerically using a Runge-Kutta scheme of order 5(4).
@@ Appendix B Response functions of the LIF neuron
The susceptibilities required in Eq. (17) are the susceptibility of the LIF neuron Eq. (4) to mean- and noise-intensity-modulations, respectively. These have been derived using Fokker-Planck theory in [27],
@@
where $\mathcal{D}_{z}(x)$ is Whittaker’s parabolic cylinder function. For the drift coefficient we need to evaluate Eqs. (54) and (55) at imaginary frequencies. This is convenient, as it only requires to evaluate $\mathcal{D}_{z}(x)$ at $z\in\mathbb{R}$. For the figures, we use the implementation $\mathtt{scipy.special.pbdv(z,x)}$.
@@ Appendix C Approximation of $D^{(2)}$
$\left\langle\dot{w}(t^{\prime})\dot{w}(t^{\prime\prime})\right\rangle_{w_{\text{traj}}(t)=w}$ $=\nu r(w)\,\delta(t^{\prime}-t^{\prime\prime})\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)^{2}$ (61)
@@
$D^{(2)}(w)$ $=\frac{1}{2}r(w)\nu\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)^{2}$
$=\frac{1}{4}r(w)\nu(\Delta_{c}^{2}\tau_{c}+r_{ac}^{2}w^{2}\tau_{ac}).$ (62)
@@ Appendix D Mean-field theory of the recurrent network
Thus, the input to all neurons is determined by the firing rate $r_{1}$, $r_{0}$, and $r_{I}$, which in turn are determined by Eq. (13). This self-consistent set of equations is solved using a damped saddle-point